OpenStudy (anonymous):

Given the equation y - 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in standard form. A. x - 2y = -12 B. y = 1/2x C. y = 1/2x + 6 D. y = 1/2x + 9

OpenStudy (anonymous):

@KamiBug one more pweeeaaassseee

OpenStudy (anonymous):

y - 3 = 1/2(x + 6) -- distribute through the parenthesis y - 3 = 1/2x + 3 -- add 3 to both sides y = 1/2x + 3 + 3 y = 1/2x + 12 --- multiply everything by 2 2y = x + 12 -- subtract x from both sides -x + 2y = 12 -- multiply by -1 to make x positive x - 2y = -12 <== standard form

OpenStudy (anonymous):

well...I feel stupid...standard form is Ax + By = C...there is only one answer in this form

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

yep..its A....you wouldn't have even had to work it out...lol

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure thing :)